OpenStudy (anonymous):

value of the function for the given x-value f(x) = 3 • 16x for x = - 3/2

OpenStudy (aravindg):

Hi there! Welcome to Openstudy :) If you have any doubts on how the site works ,feel free to ask me or any other Ambassador with purple A near them .Have great time here . Check out our COC : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Just put x = -3/2 in place of x: \[f(x) = 3 \cdot 16 x \implies f(\frac{-3}{2}) = 3 \times 16 \times \frac{-3}{2} = ??\]